Home‎ > ‎Crew 160‎ > ‎

0105-DailySummaryReport

Greetings from Crew 160! Below is our Daily Summary Report. I will be sending EVA and media photos/videos as well!

MDRS Daily Summary Report for sol 10 1/5/2016

Lost

Executive Officer Reporting

Mission Status: Today the crew worked on induvial research alongside waiting for a planned EVA to scout out further territory near White Rock Reservoir and take pictures.

Sol Activity Summary: The crew woke up ready to go for a planned EVA today to the White Rock Reservoir. We prepped ourselves with music from Star Wars, Indiana Jones, and Jurassic Park. Four of our crew hopped onto the ATVS and headed out go scout, though we didn’t get to the place at first and had to come back to the Hab to fix the situation and head back to scout again. The second team of three of the original four who went at first, went back on the second EVA and found White Rock Reservoir and took stunning photos and videos.

Look Ahead Plan: Crew is still working hard on their research and awaiting to hear about future EVA’s.

Anomalies in work: There are no anomalies to report.

Weather: The Martian sky was cloudy but little snow flurries.

Crew Physical Status: Each crew member is doing fine!

EVA: Today’s EVA was planned for four crew members to scout and take photos of White Rock Reservoir.

Reports to be filed:

Commanders Report

Journalist Report

HSO Report

GreenHab Report

EVA #09 Summary Report

EVA #10 Summary Report

Engineering Report

Support Requested: N/A